# The spectral continuum in the Rabi-Stark model

### v1

@@ 4 Numerical results for large-dimensional state space
The state space is still restricted to 200 Fock states, a truncation which is not sufficient to find all states below $E=30$. The three isolated states above the region of dense levels are preBICs (corresponding to the numbers $n=10,12,14$ in Equation (46)) and obtained in the numerical calculation because the true eigenstates with a preBIC energy have low photon content, i.e. they are located in a state space with low dimension, approximating the full Fock space $\ell^{2}(\rm{I\!N})$. But with $N_{trun}=200$ one cannot reach the states in the preContinuum which hybridize with the preBICs in Figure 2 with numbers $10$ to $14$, therefore those preBIC level lines are continuous and show no avoided crossings, in contrast to the preBICs 00 to $8$ which exhibit avoided crossings with the low-lying states of the preContinuum which is cut off at $E\sim 20$ due to $N_{trun}=200$. Remarkably, the preBIC for $n=0$ is almost perfectly approximated by the BIC for $n=0$. The hybridization between preBIC and preContinuum is shown for $n=0$ in Figure 3. The avoided crossings are already very small.
@@
For $\Delta=0.7$, there exists no lower discrete spectrum at $\gamma=1$. So we do not expect to see the lower bound states (LBS) predicted by the first confluence process in this case. The ground state should eventually reach the threshold value $E=-\Delta-2g^{2}$ for $\gamma\rightarrow 1_{-}$, without spectral gap to the excited states which form a continuum for both confluence processes. The case $\gamma=0.99$ is shown in Figure 4. While the first preBIC is well approximated by the BIC with $n=0$ for $\gamma\rightarrow 1_{-}$, the ground state is separated by a gap from $E_{thr}$, which is bigger than the average level distance below $E=-\Delta$, the upper limit of the “small” continuum predicted by the first confluence process. The difference in the density of states above and below the line $E=-\Delta$ could be interpreted in terms of an “excited state” quantum phase transition [63] for $\gamma<\omega$, as suggested by the slow-mode approximation (see section 5 below).
@@
The state space dimension is here $N_{trun}=10000$. The area $-\Delta-2g^{2}<E<-\Delta$ marks the boundaries of the “small” continuum according to the first confluence process. We see that the density of states shows no apparent change at $E=-\Delta$, so the excited state QPT is less pronounced for $\gamma$ very close to the critical point. This entails that the second confluence process predicting the “big” continuum for $E>E_{thr}$ in the limit $\gamma\rightarrow 1_{-}$ is the correct description also for the preContinuum at $\gamma=0.999$ while the “small” continuum seems to have no influence on the density of states for $\gamma\rightarrow 1_{-}$. This question will be the subject of future study. Our results show furthermore that the set of lower bound states below threshold predicted by the first confluence process is only partially present in form of precursors even for $\gamma$ rather close to $\omega$. Only one such precursor (the ground state) is visible in Figure 5.
5 Analytical results in the slow-mode approximation
A more intuitive physical picture of how the BICs and the continuum arise may be obtained by examining an approximation that leads to a pair of effective potential bands for the coupled quantum system. This approach is adapted from a technique employed for the QRM in [64] and analysed in more detail in [65]. Equation (3) may be written in terms of the position and momentum operators $\hat{q}=(2m\omega)^{-1/2}(a^{\dagger}+a)$ and $\hat{p}=i(m\omega/2)^{1/2}(a^{\dagger}-a)$ as
@@ 5 Analytical results in the slow-mode approximation
$|\psi\rangle=\psi_{+}(q)|+x\rangle+\psi_{-}(q)|-x\rangle,$ (61)
where $\sigma_{x}|\pm x\rangle=\pm|\pm x\rangle$. With $\hat{q}=q$ and $\hat{p}=-id/dq$ in this basis, the Schrödinger equations for the |±$|\pm x\rangle$ components become
$\frac{m\omega^{2}}{2}\left(1\pm\frac{\gamma}{\omega}\right)q^{2}\psi_{\pm}(q)-\frac{1}{2m}\left(1\pm\frac{\gamma}{\omega}\right)\frac{d^{2}}{dq^{2}}\pm\Delta^{\prime}\psi_{\pm}(q)+g\sqrt{2m\omega}q\,\psi_{\mp}(q)$ (62)
$=E^{\prime}\psi_{\pm}(q),$ (63)
@@
$E_{n}^{a}=\Delta^{\prime}+\left(n+\frac{1}{2}\right)\omega_{+}\left(1+\frac{2g^{2}}{\omega_{+}\Delta^{\prime}}\right)^{1/2}-\frac{\omega}{2}$ (72)
@@
$E_{b}(q)$ $=\Delta^{\prime}-\frac{1}{2}\omega+\frac{m\omega^{2}}{2}\left(1-\frac{\gamma}{\omega}-\frac{2g^{2}}{\omega\Delta^{\prime}}\right)q^{2}$ (73)
$+\left(\frac{m\omega^{2}}{2}\right)^{2}\left(\frac{2g^{4}}{\omega^{2}(\Delta^{\prime})^{3}}+\frac{2g^{2}\gamma}{\omega^{2}(\Delta^{\prime})^{2}}\right)q^{4}+\mathcal{O}(q^{6}).$
@@
$E_{n}^{b}=-\Delta^{\prime}+\left(n+\frac{1}{2}\right)\omega_{-}\left(1-\frac{2g^{2}}{\omega_{-}\Delta^{\prime}}\right)^{1/2}-\frac{\omega}{2},$ (74)
where $n=0,1,2,\dots$ and $\omega_{-}=\omega(1-\gamma/\omega)$. Formula (74) makes sense for $g$ below the value given in (75) and leads there to descending level lines. Both behaviours are displayed in Figure (7). This allows to interpret the ascending lines, which become preBICs for $\gamma\rightarrow\omega$, with the upper band (65) and the descending lines, forming the preContinuum, with the lower band (66).
Figure 7: In this spectral graph, both positive and negative parity of the numerical spectra are presented. In the meanwhile, both $E_{n}^{a}$ (green) and $E_{n}^{b}$ (blue) according to Equations (72,74) with $n=0,1,2,...,7$ are plotted.
@@
The validity of the picture for large $\Delta/\omega$ is supported by examining the energy spectra of the QRM [65]. As $\Delta/\omega$ becomes larger, the avoided crossings in the spectrum become narrower and sharper. To the eye, two sets of apparently continuous curves emerge, one curving upwards and the other downwards as a function of $g$. Although the harmonic approximation does not fully capture this behavior, especially for the lower band where the range of $g$ values for which it holds is restricted by the change from a single well to a double well, the validity of the overall picture and of the harmonic approximation both improve as $\Delta$ increases.
@@
The intuitive interpretation provided by the effective potential picture is consistent with the results in sections 3 and 4. The preBICs/BICs are identified with the bound states of the upper potential band. As $\gamma$ increases toward $\omega$, the lower band becomes wider and flatter, compressing the energy levels into a quasicontinuum. As long as $\gamma$ is less than $\omega$, the potential becomes infinite at $q=\pm\infty$, so the spectrum remains discrete and the eigenstates are normalizable. In the limit $\gamma\rightarrow\omega$, two things happen in parallel: The effective potential $E_{b}(q)$ of the lower band (66) becomes flat for $q\rightarrow\pm\infty$ and the effective mass $m/(1-\gamma/\omega)$ becomes infinite (see Equation (70)), freezing the particle at an arbitrary given position. The latter reasoning produces precisely the “small” continuum with energy values $E_{thr}<E<-\Delta$ (note that max$(E_{b}(q))=E_{b}(0)=-\Delta$ for $\gamma=\omega$), while the asymptotically flat band corresponds to scattering states with arbitrary high energy as long as the kinetic term cannot be neglected which, however, pushes these states out of the Fock space, analogous to the phenomenon associated with the second confluence process described in section 3.
Finally, for $\gamma<\omega$, the double-well structure of $E_{b}(q)$ indicates the presence of an excited state QPT with concomitant change of the density of states at $E=-\Delta$ as seen in Figure 4 which will be the subject of future investigations.
6 Conclusions
@@
We have re-examined the Rabi-Stark model using a scale transformation to an effective quantum Rabi model. The resulting $G$-function has a pole structure indicating for $\gamma\rightarrow\omega_{-}$ the emergence of a “big” continuum above the threshold energy $E_{c}=-\Delta-g^{2}/\omega$, extending to positive infinity on the real energy axis. Previous studies of the isolated point $\gamma=\omega$ [35, 38, 66] do not yield this continuum but instead a set of discrete energy levels separated by a “small” continuum for energies $-\Delta>E>E_{thr}=-\Delta-2g^{2}/\omega$. We have shown analytically that this result is associated to a confluence process of the Schrödinger equation which postulates a bounded photon content for all eigenstates. Allowing for unbounded photon content even for states with low energy leads to another confluence process and a concomitant representation of the canonical commutation relations (54) which is not unitarily equivalent to the standard one in the limit $\gamma\rightarrow\omega_{-}$, a phenomenon well-known in quantum field theory [62]. This alternative confluence process leads to the “big” continuum, as suggested by the $G$-function, but with the lower threshold energy $E_{thr}=-\Delta-2g^{2}/\omega$. Because the first confluence is still valid, we obtain from it not only the “small” as part of the “big” continuum, but also two sets of discrete states, the LBS which are true bound states below $E_{thr}$ but exist only in a limited parameter regime, and the BICs above $E=-\Delta$ which are embedded into the “big” continuum. This situation bears a remarkable resemblance to the original atomic Stark effect describing the hydrogen atom in a static and spatially constant electric field which features also “bound states” which hybridize with the surrounding continuum and are thus only metastable [67, 68, 69].
@@
We have also investigated the model with a different analytical approximation based on effective potentials which should be valid for large $\Delta/\omega$ but turns out to give excellent qualitative insight for arbitrary parameters. Especially, it is possible to identify the preBICs with the eigenstates of the upper potential surface and the preContinuum with eigenstates of the lower one. The threshold energy for the single continuum appearing here in the limit $\gamma\rightarrow\omega_{-}$, comprising both the “big” and the “small” from section 3, is the same as obtained by the other methods.
The QRSM is well-defined for all parameter values, including the regime $|\gamma|>\omega$, because the Hamilton operator is always self-adjoint, although it has no ground state for $|\gamma|>\omega$, similar to the two-photon QRM [17]. Therefore it can be studied without the regularization by an additional confining potential proposed in [66]. The central feature of the vicinity of the critical point, energetically low-lying states with very high photon content hybridizing with states having low photon content, deserves further study both theoretically and experimentally, the latter in view of possible applications for quantum technologies. Exact diagonalization of the QRSM using state spaces with high truncation number should allow to study the preContinuum and the gradual transition from normalizable states in $\ell^{2}(\rm{I\!N})$ to states outside of the Hilbert space, forming eventually the spectral continuum at the critical point. On the other hand, the slow-mode approximation lends itself to further investigation of the possible excited state quantum phase transition observed numerically for $\gamma$ not too close to $\omega$.

### v2

@@ IV Numerical results for large-dimensional state space
The state space is still restricted to 200 Fock states, a truncation which is not sufficient to find all states below $E=30$. The three isolated states above the region of dense levels are preBICs (corresponding to the numbers $n=10,12,14$ in Equation (46)) and obtained in the numerical calculation because the true eigenstates with a preBIC energy have low photon content, i.e. they are located in a state space with low dimension, approximating the full Fock space $l^{2}(\mathbb{R})$. But with $N_{trun}=200$ one cannot reach the states in the preContinuum which hybridize with the preBICs in Figure 2 with numbers $10$ to $14$, therefore those preBIC level lines are continuous and show no avoided crossings, in contrast to the preBICs 00 to $8$ which exhibit avoided crossings with the low-lying states of the preContinuum which is cut off at $E\sim 20$ due to $N_{trun}=200$. Remarkably, the preBIC $|\psi_{0}\rangle$ for $n=0$ is almost perfectly approximated by the BIC for $n=0$. This can be understood from the fact that this state has photon content $\langle\psi_{0}|a^{\dagger}a|\psi_{0}\rangle\sim 0$, as well as the corresponding BIC. Because the Hamiltonian obtained from the first confluence process in Equation (26) is equivalent to set $\gamma=\omega$ in Equations (9, 10) which removes the harmonic oscillator term in the lower spin sector, the difference between the BIC Hamiltonian in (26) and the true one for $\gamma=0.9\omega$ is very small if the photon content of the considered eigenstates is almost zero. For the same reason, the eigenvalues of BICs and preBICs deviate more strongly for higher photon content, as seen in Figure 2.
The hybridization between preBIC and preContinuum is shown for $n=0$ in Figure 3. The avoided crossings are already very small.
@@
For $\Delta=0.7$, there exists no lower discrete spectrum at $\gamma=1$. So we do not expect to see the lower bound states (LBS) predicted by the first confluence process in this case. The ground state should eventually reach the threshold value $E=-\Delta-2g^{2}$ for $\gamma\rightarrow 1_{-}$, without spectral gap to the excited states which form a continuum for both confluence processes. The case $\gamma=0.99$ is shown in Figure 4. While the first preBIC is well approximated by the BIC with $n=0$ for $\gamma\rightarrow 1_{-}$, the ground state is separated by a gap from $E_{thr}$, which is bigger than the average level distance below $E=-\Delta$, the upper limit of the “small” continuum predicted by the first confluence process.
@@
The state space dimension is here $N_{trun}=10000$. The area $-\Delta-2g^{2}<E<-\Delta$ marks the boundaries of the “small” continuum according to the first confluence process. We see that the density of states shows no apparent change at $E=-\Delta$. This entails that the second confluence process predicting the “big” continuum for $E>E_{thr}$ in the limit $\gamma\rightarrow 1_{-}$ is the correct description also for the preContinuum at $\gamma=0.999$ while the “small” continuum seems to have no influence on the density of states for $\gamma\rightarrow 1_{-}$. This question will be the subject of future study. Our results show furthermore that the set of lower bound states below threshold predicted by the first confluence process is only partially present in form of precursors even for $\gamma$ rather close to $\omega$. Only one such precursor (the ground state) is visible in Figure 5.
V Analytical results in the slow-mode approximation
A more intuitive physical picture of how the BICs and the continuum arise may be obtained by examining an approximation that leads to a pair of effective potential bands for the coupled quantum system. This approach is adapted from a technique employed for the QRM in Graham1984 and analysed in more detail in EKIthesis . Equation (3) may be written in terms of the position and momentum operators $\hat{q}=(2m\omega)^{-1/2}(a^{\dagger}+a)$ and $\hat{p}=i(m\omega/2)^{1/2}(a^{\dagger}-a)$ as
@@ V Analytical results in the slow-mode approximation
$\lvert\nobreak\psi\nobreak\rangle=\psi_{+}(q)\lvert\nobreak+x\nobreak\rangle+\psi_{-}(q)\lvert\nobreak-x\nobreak\rangle,$ (61)
where $\sigma_{x}\lvert\nobreak\pm x\nobreak\rangle=\pm\lvert\nobreak\pm x\nobreak\rangle$. With $\hat{q}=q$ and $\hat{p}=-id/dq$ in this basis, the Schrödinger equations for the |±$\lvert\nobreak\pm x\nobreak\rangle$ components become
$\frac{m\omega^{2}}{2}\left(1\pm\frac{\gamma}{\omega}\right)q^{2}\psi_{\pm}(q)-\frac{1}{2m}\left(1\pm\frac{\gamma}{\omega}\right)\frac{d^{2}}{dq^{2}}\psi_{\pm}(q)\pm\Delta^{\prime}\psi_{\pm}(q)+g\sqrt{2m\omega}q\,\psi_{\mp}(q)=E^{\prime}\psi_{\pm}(q),$ (62)
@@
$E_{n}^{a}=\Delta^{\prime}+(n+\tfrac{1}{2})\omega_{+}\left(1+\frac{2g^{2}}{\omega_{+}\Delta^{\prime}}\right)^{1/2}-\frac{\omega}{2}$ (71)
@@
$E_{b}(q)=\Delta^{\prime}-\frac{1}{2}\omega+\frac{m\omega^{2}}{2}\left(1-\frac{\gamma}{\omega}-\frac{2g^{2}}{\omega\Delta^{\prime}}\right)q^{2}+\left(\frac{m\omega^{2}}{2}\right)^{2}\left(\frac{2g^{4}}{\omega^{2}(\Delta^{\prime})^{3}}+\frac{2g^{2}\gamma}{\omega^{2}(\Delta^{\prime})^{2}}\right)q^{4}+\mathcal{O}(q^{6}).$ (72)
@@
$E_{n}^{b}=-\Delta^{\prime}+(n+\tfrac{1}{2})\omega_{-}\left(1-\frac{2g^{2}}{\omega_{-}\Delta^{\prime}}\right)^{1/2}-\frac{\omega}{2},$ (73)
where $n=0,1,2,\dots$ and $\omega_{-}=\omega(1-\gamma/\omega)$. Formula (73) makes sense for $g$ below the value given in (74) and leads there to descending level lines as function of $g$. Both behaviors are displayed in Figure (7). This allows to interpret the ascending lines, which become preBICs for $\gamma\rightarrow\omega$, with the upper band (64) and the descending lines, forming the preContinuum, with the lower band (65).
Figure 7: In this spectral graph, both positive and negative parity of the numerical spectra are presented for $\Delta=4$ and $\gamma=0.2$. In the meanwhile, both $E_{n}^{a}$ (green) and $E_{n}^{b}$ (blue) according to Equations (71,73) with $n=0,1,2,...,7$ are plotted.
@@
The validity of the picture for large $\Delta/\omega$ is supported by examining the energy spectra of the QRM EKIthesis . As $\Delta/\omega$ becomes larger, the avoided crossings in the spectrum become narrower and sharper. To the eye, two sets of apparently continuous curves emerge, one curving upwards and the other downwards as a function of $g$. Although the harmonic approximation does not fully capture this behavior, especially for the lower band where the range of $g$ values for which it holds is restricted by the change from a single well to a double well, the validity of the overall picture and of the harmonic approximation both improve as $\Delta$ increases.
@@
The intuitive interpretation provided by the effective potential picture is consistent with the results in sections III and IV. The preBICs/BICs are identified with the bound states of the upper potential band. As $\gamma$ increases toward $\omega$, the lower band becomes wider and flatter, compressing the energy levels into a quasicontinuum. As long as $\gamma$ is less than $\omega$, the potential becomes infinite at $q=\pm\infty$, so the spectrum remains discrete and the eigenstates are normalizable. In the limit $\gamma\rightarrow\omega$, two things happen in parallel: The effective potential $E_{b}(q)$ of the lower band (65) becomes flat for $q\rightarrow\pm\infty$ and the effective mass $m/(1-\gamma/\omega)$ becomes infinite (see Equation (69)), freezing the particle at an arbitrary given position. The latter reasoning produces precisely the “small” continuum with energy values $E_{thr}<E<-\Delta$ (note that max$(E_{b}(q))=E_{b}(0)=-\Delta$ for $\gamma=\omega$), while the asymptotically flat band corresponds to scattering states with arbitrary high energy as long as the kinetic term cannot be neglected which, however, pushes these states out of the Fock space, analogous to the phenomenon associated with the second confluence process described in section III.
VI Conclusions
@@
We have re-examined the Rabi-Stark model using a scale transformation to an effective quantum Rabi model. The resulting $G$-function has a pole structure indicating for $\gamma\rightarrow\omega_{-}$ the emergence of a “big” continuum above the threshold energy $E_{c}=-\Delta-g^{2}/\omega$, extending to positive infinity on the real energy axis. Previous studies of the isolated point $\gamma=\omega$ maciejewski2015 ; xie2019 ; li2023 do not yield this continuum but instead a set of discrete energy levels separated by a “small” continuum for energies $-\Delta>E>E_{thr}=-\Delta-2g^{2}/\omega$. We have shown analytically that this result is associated to a confluence process of the Schrödinger equation which postulates a bounded photon content for all eigenstates. Allowing for unbounded photon content even for states with low energy leads to another confluence process and a concomitant representation of the canonical commutation relations (54) which is not unitarily equivalent to the standard one in the limit $\gamma\rightarrow\omega_{-}$, a phenomenon well-known in quantum field theory segal1970 . This alternative confluence process leads to the “big” continuum, as suggested by the $G$-function, but with the lower threshold energy $E_{thr}=-\Delta-2g^{2}/\omega$. Because the first confluence is still valid, we obtain from it not only the “small” as part of the “big” continuum, but also two sets of discrete states, the LBS which are true bound states below $E_{thr}$ but exist only in a limited parameter regime, and the BICs above $E=-\Delta$ which are embedded into the “big” continuum. This situation bears a remarkable resemblance to the original atomic Stark effect describing the hydrogen atom in a static and spatially constant electric field which features also “bound states” which hybridize with the surrounding continuum and are thus only metastable hehenberger1974 ; cerjan1978 ; glushkov1993 .
@@
We have also investigated the model with a different analytical approximation based on effective potentials which should be valid for large $\Delta/\omega$ but turns out to give excellent qualitative insight for arbitrary parameters. Especially, it is possible to identify the preBICs with the eigenstates of the upper potential surface and the preContinuum with eigenstates of the lower one. The threshold energy for the single continuum appearing here in the limit $\gamma\rightarrow\omega_{-}$, comprising both the “big” and the “small” from section III, is the same as obtained by the other methods.
The QRSM is well-defined for all parameter values, including the regime $|\gamma|>\omega$, because the Hamilton operator is always self-adjoint, although it has no ground state for $|\gamma|>\omega$, similar to the two-photon QRM braak2023 . Therefore it can be studied without the regularization by an additional confining potential proposed in li2023 . The central feature of the vicinity of the critical point, energetically low-lying states with very high photon content hybridizing with states having low photon content, deserves further study both theoretically and experimentally, the latter in view of possible applications for quantum technologies. Exact diagonalization of the QRSM using state spaces with high truncation number should allow to study the preContinuum and the gradual transition from normalizable states in $l^{2}(\mathbb{R})$ to states outside of the Hilbert space, forming eventually the spectral continuum at the critical point.
Acknowledgements.
@@
Corresponding author. Email: daniel.braak@uni-a.de
∗∗ Corresponding author. Email: congllzu@gmail.com